A STONE IS THROWN STRAIGHT UP INTO THE AIR AND HAS A HANG TIME OF 5 SECONDS FIND THE HEIGHT TO WHICH THE BALL REACHES ITS PEAK


1
Expert's answer
2021-03-24T19:41:28-0400

Answer

Using newton second equation of motion

"v=u+at"

"0=u-gt"

u=gt

So maximum height

"H=\frac{u^2}{2g}\\=\frac{g^2t^2}{2g}\\=\frac{gt^2}{2}\\=\frac{10\times5\times5}{2}\\=125m"